A block B of mass 5kg is placed on another block A of mass 10kg, which rests on a smooth horizontal surface. If u=0.4 between A and B and a force F=40N is applied on block B, the acceleration of A is :
​a) 3m/s2
​b) 2 m/s2
​c) 4 m/s2
​d) 8/3 m/s2

Grade:11

1 Answers

Shivam Agrawal
4 years ago
Assume the force on Block B to be applied in right Direction. Let that Force be F1= 40N 
Frictional Force acting between the two block will act in the opposite direction i.e is left direction. Let this force be F2.
F2= u* (mass of B) * g
    = 0.4*5*10
F2 = 20N 
This force will create acceleration in Block A. 
We know Force = Mass * Acceleration 
F2 = (mass of A)* (Acceleration of A)
20 = 10 *( Acceleration of A)
Acceleration of A = 2m/s2.